OpenStudy (jamaljunior):

5+ 2(3x - 2) = 13 @ganeshie8

OpenStudy (mikey):

5+6x-4=13

OpenStudy (jamaljunior):

what do i do?

OpenStudy (mikey):

try to solve for x

OpenStudy (mikey):

so what would u do first

OpenStudy (jamaljunior):

subtract from both sides to seperate x

OpenStudy (jamaljunior):

yes or no?

OpenStudy (mikey):

correct

OpenStudy (jamaljunior):

can you tell me number to subract?

OpenStudy (mikey):

add 4 then subtract 5

OpenStudy (jamaljunior):

add 4 to what?

OpenStudy (mikey):

13

OpenStudy (jamaljunior):

so 6x - 5 = 1 and 17 - 5 = 12?

OpenStudy (mikey):

first part is wrong u only add and subtract like terms on the opposite side of the = sign

OpenStudy (mikey):

so it would be 6x=12

OpenStudy (jamaljunior):

then 6/6 and 12/6?

OpenStudy (mikey):

correct so x=2

OpenStudy (jamaljunior):

thanks @mikey